Assume that in January 2017, the average house price in a particular area was $310,450. In January 2001, the average price was $204,300.

What was the annual increase in selling price? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Annual increase in selling price _____ %
Answer #1

Annual increase in selling price =(Selling Price in 2017/Selling Price in 2001)^(1/n)-1 =(310450/204300)^(1/16)-1 = 2.65%